OpenStudy (anonymous):

A snowmobile has an initial velocity of 3.0 m/s Part 1: If it accelerates at the rate of 0.5 m/s/s for 7.0 sec, what is the final vel.? Part 2: If it accelerates at the rate of -0.60m/s/s, how long will it take to reach a complete stop? **for part 1, is this how you do it? final velocity=3+0.5*7=3+3.5=6.5 m/s ? and not sure at all for part 2! :/ @phi @Michele_Laino what do y'all think? :) thank you!!

OpenStudy (michele_laino):

please, note that for part 1, you have to apply the same formula used before, namely \[v(t)=v _{0}+a*t\]

OpenStudy (anonymous):

yes:) did i do that right? v0=3.0 + 0.5 * 7.0 = 3.0 + 3.5 = 6.5 m/s final velocity?

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

for part 2, you have to apply this formula: \[0=v _{0}-a*t\] where v_0=6.5 and a=0.6, so we have: \[0=6.5-0.6*t\] please solve for t

OpenStudy (anonymous):

ah okay, so then we get 0=5.9*t t= - 5.9 ?

OpenStudy (anonymous):

it will be negative?

OpenStudy (michele_laino):

no, please note that t is time!

OpenStudy (anonymous):

oh! so t will be 5.9 seconds?

OpenStudy (michele_laino):

sorry, my equation is: \[0=6.5-(0.6*t)\]

OpenStudy (anonymous):

so it will be 0=6.5-0.6t 0.6t=6.5 t=10.83 ?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

oh yay!! thank you so much!! :)

OpenStudy (michele_laino):

thank you! :)